# VisRet: Visualization Improves Knowledge-Intensive Text-to-Image Retrieval

### v2

@@ Abstract
Text-to-image retrieval (T2I retrieval) remains challenging because cross-modal embeddings often behave as bags of concepts and underrepresent structured visual relationships such as pose and viewpoint. We propose Visualize-then-Retrieve (VisRet), a new paradigm for T2I retrieval that mitigates this limitation of cross-modal similarity alignment. VisRet first projects textual queries into the image modality via T2I generation. Then, it performs retrieval within the image modality to bypass the weaknesses of cross-modal retrievers in recognizing subtle visual-spatial features. Across four benchmarks (Visual-RAG, INQUIRE-Rerank, Microsoft COCO, and our new Visual-RAG-ME featuring multi-entity comparisons), VisRet substantially outperforms cross-modal similarity matching and baselines that recast T2I retrieval as text-to-text similarity matching, improving nDCG@30 by 0.125 on average with CLIP as the retriever and by 0.121 with E5-V. For downstream question answering, VisRet increases accuracy on Visual-RAG and Visual-RAG-ME by 3.8% and 15.7% in top-1 retrieval, and by 3.9% and 11.1% in top-10 retrieval. Ablation studies show compatibility with different T2I instruction LLMs, T2I generation models, and downstream LLMs. VisRet provides a practical and principled path that energizes further advances in vision-language retrieval. Our code and the new benchmark are publicly available at https://github.com/xiaowu0162/Visualize-then-Retrieve.
@@ 1 Introduction
Text-to-Image (T2I) retrieval is the task of selecting the most relevant images from a visual corpus based on a textual query. It plays a crucial role in enabling knowledge-intensive applications that require supporting textual inputs with rich visual content (Chen et al., 2022; Wang et al., 2023; Sheynin et al., 2023; Braun et al., 2024).
A common approach to T2I retrieval is to embed both the text query and image candidates into a shared representation space, where similarity scores are computed (Frome et al., 2013; Kiros et al., 2014). However, obtaining accurate similarity rankings that capture fine-grained semantic alignments between text and image remains a long-standing challenge. Prior studies have observed that cross-modal embeddings often behave like “bags-of-concepts” for similarity matching, failing to model structured relationships among visual elements (Yüksekgönül et al., 2023; Kamath et al., 2023). For instance, Figure˜1 presents a query that requires images of an entity (a Barnacle Goose) at specific postures (wings unfolded) to answer. While the embedding model succeeds at matching the entity type, it struggles to recognize subtler visual-spatial features such as the pose of the wing (unfolded) and the camera perspective (upward-facing shot). To address these limitations, existing work has explored improving the embedding quality (Radford et al., 2021; Yu et al., 2022), query reformulation (Levy et al., 2023), and multi-stage reranking pipelines (Liu et al., 2024; Feng et al., 2025). Nevertheless, all these strategies are ultimately constrained by the intrinsic difficulty of cross-modal similarity alignment, as they cannot bypass the stage of text-to-image similarity search.
@@
We propose Visualize-then-Retrieve (VisRet), a novel retrieval paradigm that decomposes T2I retrieval into two stages: text-to-image modality projection followed by within-modality retrieval. Concretely, VisRet first utilizes a text-to-image generation (T2I generation) model to visualize the textual query, highlighting the key desired visual-spatial features. Then, the visualized query is used to perform direct image-to-image retrieval.
Compared to prior methods, VisRet offers two key advantages. First, visualizations provide a more expressive and intuitive medium for representing multiple compositional concepts, including entities, poses, and spatial relations. These concepts are often difficult to convey through text alone, and attempting to encode them exhaustively in a text query can impair cross-modal matching performance due to limitations in embedding quality. In contrast, as shown in Figure˜1, the visualized query is able to accurately depict the desired entity, posture, and camera angle at the same time. Second, by operating entirely within the image modality during retrieval, VisRet avoids the weaknesses of cross-modal retrievers and instead leverages the stronger capacity of these retrievers in unimodal retrieval (Koishigarina et al., 2025).
To investigate the effectiveness of VisRet, we conduct exhaustive experiments on two commonly used cross-modal embedding models using four benchmarks: Visual-RAG (Wu et al., 2025b), INQUIRE-Rerank (Vendrow et al., 2024), Microsoft COCO (Chen et al., 2015), and Visual-RAG-ME, a new benchmark we introduce that features challenging visual feature comparison questions across multiple entities. Results show that VisRet substantially outperforms baseline cross-modal similarity matching as well as three strong baseline methods that casts T2I retrieval to text-to-text similarity matching (section˜5.2). When CLIP (Radford et al., 2021) is used as the retriever, VisRet outperforms the baselines by 0.125 higher nDCG@30, averaged over four benchmarks. With E5-V (Jiang et al., 2024) as the retriever, the performance gain becomes 0.121. Moreover, VisRet enhances downstream performance in retrieval-augmented generation (RAG) settings (section˜5.3). It improves T2I question answering accuracy on Visual-RAG and Visual-RAG-ME by 3.8% and 15.7% in top-1 retrieval setting and 3.9% and 11.1% in top-10 retrieval setting. Further analyses show that VisRet is compatible with different T2I instruction generation and downstream question answering LLMs. In addition, the retrieval performance consistently improves with stronger T2I models. Finally, although they are effective in retrieving visual knowledge, the generated image queries themselves are generally insufficient as standalone knowledge to answer downstream questions.
@@ 2 Related Work
Early T2I retrieval benchmarks evaluate the ability to identify images based on their paired human-written captions. These datasets span multiple domains and include widely used benchmarks such as COCO (Chen et al., 2015), Flickr8K (Hodosh et al., 2013), Flickr30K (Young et al., 2014), and Fashion200K (Han et al., 2017). As multimodal embedding models have matured, more challenging benchmarks have been introduced to assess retrieval in knowledge-intensive settings. These newer datasets, such as WebQA (Chang et al., 2022), INQUIRE (Vendrow et al., 2024), Visual-RAG (Wu et al., 2025b), and MRAG-Bench (Hu et al., 2024), shift the focus from caption matching to retrieving images that contain the knowledge necessary to answer complex natural language questions. These tasks challenge T2I retrieval systems to support downstream reasoning in retrieval-augmented generation (RAG) pipelines. Further extending this stream of work, our Visual-RAG-ME benchmarks T2I retrieval for reasoning over the visual features across multiple entities.
@@
Extensive research has focused on improving text-to-image retrieval from different perspectives. First, a number of works aim to train better multimodal embeddings by designing better training objectives and data mixtures (Faghri et al., 2018; Radford et al., 2021; Yu et al., 2022; Li et al., 2022). Other studies improve various stages in the retrieval pipeline, such as textual query expansion (Levy et al., 2023; Lee et al., 2024) and reranking (Liu et al., 2024; Feng et al., 2025; Ding et al., 2025). Finally, a recent line of work introduces generative image retrieval (Li et al., 2024; Qu et al., 2025), which trains a generative model to directly memorize an index of the image corpus. Different from these existing approaches, VisRet expands the query semantics by directly visualizing it in the image space, thereby alleviating the workload of cross-modal retrieval. In addition, VisRet is a training-free plug-and-play framework that can accommodate any off-the-shelf retriever and pre-computed image embedding indices.
@@ 3.1 Problem Formulation
Given a textual query $q$ and an image corpus $\mathcal{I}$, the task of Text-to-Image retrieval aims to retrieve $n\geq 1$ images $y_{1},...,y_{n}\in\mathcal{I}$ that best correspond to the semantics of $q$. Our paper further considers the task of Visual Question Answering (VQA), where the query is a knowledge-seeking question, with an expected answer $a$.
In this paper, we consider a basic retrieval-augmented generation (RAG) VQA pipeline: A multimodal retriever $\mathcal{R}$ retrieves $k$ images from $\mathcal{I}$, denoted as $\{r_{1},...,r_{k}\}\equiv\mathcal{R}(q,\mathcal{I})\subseteq\mathcal{I}$. Then, a large vision-language model (LVLM) $\mathcal{M}$ directly generates the answer based on the question and the retrieval results $\mathcal{M}(q,\mathcal{R}(q,\mathcal{I}))$.
@@ 3.2 Visualize-then-Retrieve
We introduce Visualize-then-Retrieve (VisRet), a two-stage T2I retrieval pipeline that bridges the modality gap through modality projection. Figure˜1 illustrates the pipeline with an intuitive example.
@@
The first stage of VisRet leverages a T2I generation system $\mathcal{T}$ to directly generate $m$ visualizations $\{v_{1},...,v_{m}\}\equiv\mathcal{T}(q)$. Concretely, based on the original query, an LLM first operates in the text space to draft T2I instruction $q^{\prime}$, describing the images that likely satisfy the feature requirements implied in $q$. Next, we use an existing T2I generation method, such as Stable Diffusion (Esser et al., 2024), to project $q^{\prime}$ into a list of image queries $\{v_{1},...,v_{m}\},v_{i}\in\mathcal{I}$. To generate diverse images, randomness can be injected either into $q^{\prime}$ or into the T2I generation. In this work, we sample $m$ times with the same $q^{\prime}$.
@@
In the second stage, VisRet performs retrieval exclusively within the image modality. First, each synthesized image $v_{i}\in\{v_{1},...,v_{m}\}$ is independently used to retrieve a ranked list of images from the corpus:
$\mathcal{R}(v_{i},\mathcal{I})=[r_{1}^{(i)},...,r_{k}^{(i)}],$
Then, we apply Reciprocal Rank Fusion (RRF) (Cormack et al., 2009) to aggregate the $m$ retrieval results. RRF assigns a fusion score to each candidate image $r$ based on its rank across $m$ lists:
@@
where $\text{rank}_{i}(r)$ is the rank position of image $r$ in list $\mathcal{R}(v_{i},\mathcal{I})$, and $\lambda$ is a hyperparameter that controls the influence of lower-ranked items. The final top-$k$ retrieval result is obtained by selecting the highest-scoring images according to $\text{score}_{\text{RRF}}(r)$.
@@ 4 Visual-RAG-ME Benchmark
To facilitate the design of intelligent T2I retrieval systems, we introduce a multi-entity question answering and T2I retrieval dataset Visual-RAG-ME as an extension of Visual-RAG. Concretely, Visual-RAG-ME constructs questions that compare the visual features between the organism covered in Visual-RAG with another similar entity, for which we manually annotate a new set of retrieval labels. The Visual-RAG-ME annotation pipeline consists of three steps as described below.
@@
The goal of this step is to identify entities that are biologically close to the original entities in Visual-RAG so that plausible and challenging questions can be constructed. For this purpose, for each entity covered in Visual-RAG, we use BM25 (Robertson and Walker, 1994) to retrieve ten entities with the most similar scientific names.
@@
In this step, we (the authors) manually review all the 391 questions in Visual-RAG and attempt to construct a corresponding multi-entity question. A question is constructed when we can identify images for the second entity that clearly depict the same feature as in the positive images for the original entity in Visual-RAG. The questions we compose generally take a comparison style that asks whether the two organisms have the same feature or which organism features a more extreme stylistic feature (e.g., lighter coloration, smoother surface etc.). We were able to construct 82 multi-entity questions after this step. Then, a round of filtering is performed to remove (1) the questions that both GPT-4o and GPT-4o-mini can answer correctly without image information and (2) questions that cover overly similar topics. The final Visual-RAG-ME contains 50 high-quality multi-entity queries.
@@
For each question, we collect images of the second entity from the iNaturalist 2021 dataset (Horn et al., 2021) and annotate their retrieval labels. A positive label is assigned only if the image clearly displays the feature required to answer the question. For some of the questions, we find that a large number of positive images exist in the iNaturalist database. We therefore implement a filtering step where at most 50 positive images are kept for each entity. Table˜3 and Appendix Table˜8 show examples with their questions and ground-truth images.
In Table˜1, we present the basic statistics of Visual-RAG-ME. While it has slightly more positive per entity compared to Visual-RAG, our new benchmark is still challenging, with both CLIP and E5-V achieving a low Recall@1 due to its lengthy and knowledge-intensive queries.
@@ 5.1 Experimental Setup
We evaluate on four challenging benchmarks: (1) Visual-RAG (Wu et al., 2025b), a T2I retrieval and VQA benchmark featuring visual knowledgeintensive questions on features of natural species that are not commonly documented in text corpora. (2) Visual-RAG-ME, a new benchmark we introduce featuring queries that compare the same visual feature across multiple entities. (3) INQUIRE-Rerank (Vendrow et al., 2024), a T2I retrieval benchmark requiring accurate knowledge of species appearance and behavior. (4) Microsoft COCO 2017 (Chen et al., 2015), a widely used image captioning dataset, which we repurpose for caption-to-image retrieval evaluation. For (3) and (4), we perform additional filtering to retain the most difficult queries and call the resulting datasets INQUIRE-Rerank-Hard and COCO-Hard. Table˜1 shows the basic statistics of each dataset. We present additional details in Appendix Appendix˜A.
@@
For all the three benchmarks, we evaluate T2I retrieval with Recall@$k$ and nDCG@$k$ with $k$ = 1, 10, 30111As nDCG@1 is undefined, we use Recall@1 as its value.. For Visual-RAG and Visual-RAG-ME, we additionally use an LLM judge to evaluate the end-to-end VQA accuracy, using a prompt derived from Visual-RAG (Wu et al., 2025b).
@@
Using the benchmarks, we evaluate three families of T2I retrieval methods:
@@
    Text-to-Image Matching: Retrieving with the original textual query against the image index using off-the-shelf retrievers (CLIP, E5-V). We also compare with a baseline of applying query rewriting via an LLM to highlight the desired features and then performing standard T2I retrieval over the same image embedding index.
@@
    Text-to-Text Matching: Converting the images into textual surrogates for text-to-text retrieval. We compare with (1) captioning the entire corpus with BLIP, (2) top-k reranking with VISA, and (3) an ablated version of VISA, where we rerank top-k images using LLM-generated captions. We use Qwen2.5-70B-VL (Bai et al., 2025) as the model for (2) and (3) and rerank the top-30 retrieval.
@@
    Image-to-Image Matching: We evaluate VisRet and an image ablation of VisRet that only uses a single image instead of multiple images (-multi-image).
@@
For all the retrieval experiments, we use CLIP and E5-V as the retrievers, GPT-4o (OpenAI, 2023b) as T2I instruction generation LLM, and gpt-image-1 (OpenAI, 2025) as the T2I Model to generate $m=3$ images. For the downstream question answering (Visual-RAG and Visual-RAG-ME), we use GPT-4o as the reader LLM. We present the full implementation details of VisRet and the baselines in Appendix˜B.

### v3

@@ Abstract
Text-to-image retrieval (T2I retrieval) remains challenging because cross-modal embeddings often behave as bags of concepts, underrepresenting structured visual relationships such as pose and viewpoint. We propose Visualize-then-Retrieve (VisRet), a retrieval paradigm that mitigates this limitation of cross-modal similarity alignment. VisRet first projects textual queries into the image modality via T2I generation, then performs retrieval within the image modality to bypass the weaknesses of cross-modal retrievers in recognizing subtle visual-spatial features. Across four benchmarks (Visual-RAG, INQUIRE-Rerank, Microsoft COCO, and our new Visual-RAG-ME featuring multi-entity comparisons), VisRet substantially outperforms cross-modal similarity matching and baselines that recast T2I retrieval as text-to-text similarity matching, improving nDCG@30 by 0.125 on average with CLIP as the retriever and by 0.121 with E5-V. For downstream question answering, VisRet increases accuracy on Visual-RAG and Visual-RAG-ME by 3.8% and 15.7% in top-1 retrieval, and by 3.9% and 11.1% in top-10 retrieval. Ablation studies show compatibility with different T2I instruction LLMs, T2I generation models, and downstream LLMs. VisRet provides a simple yet effective perspective for advancing in text-image retrieval. Our code and the new benchmark are publicly available at https://github.com/xiaowu0162/Visualize-then-Retrieve.
@@ 1 Introduction
Text-to-Image (T2I) retrieval selects the most relevant images from a visual corpus given a textual query. It plays a crucial role in knowledge-intensive applications that augment textual inputs with rich visual evidence (DBLP:conf/emnlp/ChenHCVC22; DBLP:conf/acl/Wang0LLCJHXG23; DBLP:conf/iclr/SheyninAPSGNT23; DBLP:journals/corr/abs-2412-10510).
A common approach embeds both the text query and image candidates into a shared representation space and ranks candidates by similarity (frome2013devise; DBLP:journals/corr/KirosSZ14). However, producing rankings that reflect fine-grained alignments between text and image remains a long-standing challenge. Prior studies show that cross-modal embeddings often behave like “bags of concepts” for similarity matching and fail to capture structured relationships among visual elements (DBLP:conf/iclr/Yuksekgonul0KJ023; DBLP:conf/emnlp/KamathHC23). For example, Figure˜1 presents a query that requires images of an entity (a Barnacle Goose) in a specific posture (wings unfolded). While the embedding model matches the entity type, it struggles to recognize subtler visual-spatial features such as the wing pose and the camera perspective (an upward-facing shot). To address these limitations, existing work improves embedding quality (DBLP:conf/icml/RadfordKHRGASAM21; DBLP:journals/tmlr/YuWVYSW22), performs query reformulation (levy2023chatting), or applies multi-stage reranking pipelines (DBLP:journals/tmlr/0002STG24; DBLP:conf/aaai/0001B00KZGL25). Nevertheless, these strategies remain constrained by the intrinsic difficulty of cross-modal similarity alignment because they cannot bypass the text-to-image similarity search stage.
@@
We propose Visualize-then-Retrieve (VisRet), a retrieval paradigm that decomposes T2I retrieval into two stages: text-to-image modality projection followed by within-modality retrieval. VisRet first uses a text-to-image generation (T2I generation) model to visualize the textual query, with the goal of making key visual-spatial requirements explicit. The resulting visualization is then used as a query for direct image-to-image retrieval.
Compared to prior methods, VisRet offers two advantages. First, visualizations provide a more expressive and intuitive medium for representing compositional concepts such as entities, poses, and spatial relations. These requirements can be difficult to communicate through text alone, and encoding them exhaustively in a text query can hurt cross-modal matching due to limitations in embedding quality. In contrast, as shown in Figure˜1, the visualized query can depict the desired entity, posture, and camera angle at the same time. Second, by operating entirely within the image modality during retrieval, VisRet avoids the weaknesses of cross-modal retrievers and leverages the stronger capacity of these retrievers in unimodal retrieval (DBLP:journals/corr/abs-2502-03566).
To evaluate VisRet, we conduct experiments with two widely used cross-modal embedding models across four benchmarks: Visual-RAG (visual-rag), INQUIRE-Rerank (DBLP:conf/nips/VendrowPSBJABH24), Microsoft COCO (mscoco), and Visual-RAG-ME, a new benchmark we introduce that features challenging visual feature comparison questions across multiple entities. VisRet substantially outperforms baseline cross-modal similarity matching and three strong baseline methods that cast T2I retrieval as text-to-text similarity matching (section˜5.2). When CLIP (DBLP:conf/icml/RadfordKHRGASAM21) is used as the retriever, VisRet improves nDCG@30 by 0.125 on average across the four benchmarks. With E5-V (DBLP:journals/corr/abs-2407-12580) as the retriever, the improvement is 0.121. VisRet also improves downstream performance in retrieval-augmented generation (RAG) settings (section˜5.3), increasing T2I question answering accuracy on Visual-RAG and Visual-RAG-ME by 3.8% and 15.7% in top-1 retrieval and by 3.9% and 11.1% in top-10 retrieval. Ablation studies and analyses show that VisRet is robust to the choice of instruction-generation LLM, retriever, and downstream reader LVLM, while performance depends more strongly on the T2I generation model, with stronger generators yielding larger gains. We further analyze repeated runs and find that aggregating multiple visualizations is beneficial, while using a single visualization only slightly reduces overall performance. Finally, although visualized queries are effective for retrieval, they are generally insufficient as standalone knowledge to answer downstream questions. Our code and the new Visual-RAG-ME benchmark will be released publicly to facilitate future research.
@@ 2 Related Work
Early T2I retrieval benchmarks evaluate whether a system can retrieve an image given its paired human-written caption. These datasets span multiple domains and include widely used benchmarks such as COCO (mscoco), Flickr8K (DBLP:journals/jair/HodoshYH13), Flickr30K (DBLP:journals/tacl/YoungLHH14), and Fashion200K (han2017automatic). As multimodal embedding models have matured, newer benchmarks have been introduced to assess retrieval in knowledge-intensive settings. Datasets such as WebQA (DBLP:conf/cvpr/ChangCNGSB22), INQUIRE (DBLP:conf/nips/VendrowPSBJABH24), Visual-RAG (visual-rag), and MRAG-Bench (DBLP:journals/corr/abs-2410-08182) shift the focus from caption matching to retrieving images that provide the information needed to answer knowledge-intensive questions. These tasks evaluate T2I retrieval systems as components of retrieval-augmented generation (RAG) pipelines and emphasize support for downstream question answering. Extending this line of work, Visual-RAG-ME benchmarks T2I retrieval for reasoning about visual features shared across multiple entities.
@@
A large body of work improves T2I retrieval from several angles. First, many approaches aim to train stronger multimodal embeddings through improved objectives and data mixtures (DBLP:conf/bmvc/FaghriFKF18; DBLP:conf/icml/RadfordKHRGASAM21; DBLP:journals/tmlr/YuWVYSW22; DBLP:conf/icml/0001LXH22). Other studies target specific stages of the pipeline, including textual query expansion (levy2023chatting; DBLP:conf/acl/LeeYPYY24) and reranking (DBLP:journals/tmlr/0002STG24; DBLP:conf/aaai/0001B00KZGL25; ding2025visual). Finally, a recent line of work explores generative image retrieval (DBLP:conf/acl/LiWQNLC24; qu2025tigerunifyingtexttoimagegeneration), which trains a generative model to directly memorize an index of the image corpus. In contrast, VisRet expands query semantics by visualizing the query in the image space, which reduces the burden on cross-modal retrieval. It is also training-free and plug-and-play, accommodating off-the-shelf retrievers or pre-computed image embedding indices.
@@ 3.1 Problem Formulation
Given a textual query $q$ and an image corpus $\mathcal{I}$, Text-to-Image Retrieval aims to retrieve $n\geq 1$ images $y_{1},\ldots,y_{n}\in\mathcal{I}$ that best match the semantics of $q$. We further consider Visual Question Answering (VQA), where the query is a knowledge-seeking question with an expected answer $a$.
We consider a basic retrieval-augmented generation (RAG) pipeline for VQA. A multimodal retriever $\mathcal{R}$ retrieves $k$ images from $\mathcal{I}$, denoted as $\{r_{1},\ldots,r_{k}\}\equiv\mathcal{R}(q,\mathcal{I})\subseteq\mathcal{I}$. A large vision-language model (LVLM) $\mathcal{M}$ then generates an answer conditioned on the question and retrieval results, i.e., $\mathcal{M}(q,\mathcal{R}(q,\mathcal{I}))$.
@@ 3.2 Visualize-then-Retrieve
We introduce Visualize-then-Retrieve (VisRet), a two-stage T2I retrieval pipeline that bridges the modality gap through modality projection. Figure˜1 illustrates the pipeline with an example.
@@
In the first stage, VisRet uses a T2I generation system $\mathcal{T}$ to synthesize $m$ visualizations $\{v_{1},\ldots,v_{m}\}\equiv\mathcal{T}(q)$. Concretely, given the original query, an LLM first operates in the text space to draft a T2I instruction $q^{\prime}$ that describes images likely to satisfy the feature requirements implied by $q$. We then use an existing T2I generation method, such as Stable Diffusion (DBLP:conf/icml/EsserKBEMSLLSBP24), to project $q^{\prime}$ into a set of image queries $\{v_{1},\ldots,v_{m}\}$, where each $v_{i}$ is a synthesized image. To encourage diversity, randomness can be introduced either in the instruction drafting step or within the T2I generation process. In this work, we sample $m$ times using the same $q^{\prime}$.
@@
In the second stage, VisRet performs retrieval entirely within the image modality. Each synthesized image $v_{i}\in\{v_{1},\ldots,v_{m}\}$ is independently used to retrieve a ranked list of images from the corpus:
$\mathcal{R}(v_{i},\mathcal{I})=[r_{1}^{(i)},\ldots,r_{k}^{(i)}].$
We then apply Reciprocal Rank Fusion (RRF) (DBLP:conf/sigir/CormackCB09) to aggregate the $m$ ranked lists. RRF assigns each candidate image $r$ a fusion score based on its rank across the $m$ lists:
@@
where $\text{rank}_{i}(r)$ is the rank position of image $r$ in $\mathcal{R}(v_{i},\mathcal{I})$, and $\lambda$ is a hyperparameter that controls the influence of lower-ranked items. The final top-$k$ results are obtained by selecting images with the highest $\text{score}_{\text{RRF}}(r)$.
@@ 4 Visual-RAG-ME Benchmark
To support the development of stronger T2I retrieval systems, we introduce Visual-RAG-ME, a multi-entity question answering and T2I retrieval dataset that extends Visual-RAG. Visual-RAG-ME constructs questions that compare visual features between the organism in a Visual-RAG query and another similar entity, and we manually annotate a new set of retrieval labels for the second entity. The Visual-RAG-ME annotation pipeline consists of the following three steps.
@@
This step identifies potential entities that are biologically close to the original Visual-RAG entities, enabling plausible and challenging comparison questions. For each entity in Visual-RAG, we use BM25 (10.1007/978-1-4471-2099-5_24) to retrieve up to ten entities with the most similar scientific names.
@@
We (the authors) manually review all 391 questions in Visual-RAG and attempt to construct a corresponding multi-entity question. We construct a question only when we can identify images for the second entity that clearly depict the same feature as the positive images for the original entity in Visual-RAG. The questions we compose typically use a comparison format that asks whether the two organisms share a feature or which organism exhibits a more extreme stylistic feature (e.g., lighter coloration, smoother surface). This process yields 82 multi-entity questions. We then filter out (1) questions answerable by both GPT-4o and GPT-4o-mini without image information and (2) excessive questions covering the same topic. The final Visual-RAG-ME contains 50 high-quality multi-entity queries.
@@
For each question, we collect images of the second entity from the iNaturalist 2021 dataset (DBLP:conf/cvpr/HornCBWBA21) and annotate retrieval labels. We assign a positive label only when the image clearly displays the feature required to answer the question. For some questions, many positive images exist in iNaturalist, so we apply a filtering step that keeps at most 50 positive images for each entity. Table˜5 and Appendix Table˜10 show example questions and their ground-truth images.
@@
Table˜1 reports the basic statistics of Visual-RAG-ME. While it contains slightly more positives per entity than Visual-RAG, it remains challenging. Both CLIP and E5-V achieve low Recall@1 due to the lengthy and knowledge-intensive queries.
5 Main Results
@@ 5.1 Experimental Setup
We evaluate on four benchmarks: (1) Visual-RAG (visual-rag), a T2I retrieval and VQA benchmark featuring visual knowledge-intensive questions about natural species, where the relevant evidence is often not readily available in text corpora. (2) Visual-RAG-ME, our new benchmark that compares the same visual feature across multiple entities. (3) INQUIRE-Rerank (DBLP:conf/nips/VendrowPSBJABH24), which requires accurate knowledge of species appearance and behavior for retrieval. (4) Microsoft COCO 2017 (mscoco), a widely used image captioning dataset that we repurpose for caption-to-image retrieval evaluation. For (3) and (4), we further filter queries to retain the most difficult cases, yielding INQUIRE-Rerank-Hard and COCO-Hard. Table˜1 reports dataset statistics, and we provide additional details in Appendix Appendix˜A.
@@
For all four retrieval benchmarks, we report Recall@$k$ and nDCG@$k$ with $k\in\{1,10,30\}$111nDCG@1 is undefined and we use Recall@1 as its value.. For Visual-RAG and Visual-RAG-ME, we also evaluate end-to-end VQA accuracy using an LLM judge with a prompt derived from Visual-RAG (visual-rag).
@@
We compare three families of T2I retrieval methods:
@@
    Text-to-Image Matching: Retrieve using the original textual query against the image index with off-the-shelf retrievers (CLIP, E5-V). We also include a query rewriting baseline, where an LLM rewrites the query to highlight desired features before performing standard T2I retrieval over the same embedding index.
@@
    Text-to-Text Matching: Convert images into textual surrogates and perform text-to-text retrieval avoiding cross-modal similarity matching. We compare with (1) captioning the entire corpus with BLIP, (2) top-k reranking with VISA, and (3) an ablated version of VISA, where we rerank top-k images using LLM-generated captions.
@@
    Image-to-Image Matching: At the time of writing (December 2025), VisRet is the only method in this category.
@@
For all retrieval experiments, we use CLIP and E5-V as retrievers, GPT-4o (OpenAI2023GPT4TR) as the T2I instruction generation LLM, and gpt-image-1 (openai_gpt_image_1) as the T2I model to generate $m=3$ images. For downstream question answering on Visual-RAG and Visual-RAG-ME, we use GPT-4o as the reader LLM. Full implementation details for VisRet and baselines are provided in Appendix Appendix˜B.